OpenStudy (harrisonshark):

Given the system of equations, what is the solution? x + 3y = 5 x - 3y = -1 {(-4, -3)} {(2, 1)} {(-2, 3)}

OpenStudy (mathmale):

Solve this system of linear equations simultaneously: x + 3y = 5 x - 3y = -1 This means coming up with values of x and y that "satisfy" both equations. There are several methods of solution that you could use. For now, let's concentrate on solution by addition and subtraction. First off, would you please ADD the two given equations together, column by column? Type in here what you get as your result, after simplifying it.

OpenStudy (harrisonshark):

-2 , -3??

OpenStudy (harrisonshark):

3 not -3

OpenStudy (mathmale):

Harrison, why not just do as I asked and combine the 2 equations? x + 3y = 5 x - 3y = -1 ---------- Add: x+x. What do you get? write the sum in the first column under the dashed line. combine 3y-3y. What do you get? write the result in the 2nd col. under the d. line. 5-1+ ? Write your result under the 3rd column, under the dashed line.

OpenStudy (mathmale):

I would like to see your work, and would to see it labelled. Otherwise, how am I (or how is anyone else) supposed to know what you mean by -2, -3?

OpenStudy (mathmale):

Harrison? x + 3y = 5 x - 3y = -1 ---------- 2x = 4 so x=2. Substituting x=2 into the first equation, above, results in what new equation? Solve it for y, please.